# Zero-Knowledge Federated Learning: A New Trustworthy and Privacy-Preserving Distributed Learning Paradigm

### v1

@@ IV-A Threat and Trust Model
  • Malicious clients can engage in poisoning attacks by submitting unreliable models with poor performance, aiming to disrupt the overall learning process.
@@ IV-B Veri-CS-FL
    Each client $C_{k}$ trains an updated local model $w_{k}^{t}$ from the last round global model $w^{t-1}$ utilizing its local dataset $D_{k}$, $k\in{1,2,\cdots,K}$. Subsequently, each client computes a performance metric of the updated local model $w_{k}^{t}$ by comparing the updated local model $w_{k}^{t}$ with the benchmark model $w_{s}^{t}$, and employs the proof generation algorithm of zk-SNRAK to generate a zero-knowledge proof $\pi_{k}$ for the performance metric of $w_{k}^{t}$: $\pi_{k}\leftarrow PROVE(PK,x,W)$, where the public input $x$ is the benchmark model of the server $w_{s}^{t}$ together with the hash of the updated local model $h_{k}^{t}=\text{hash}(w_{k}^{t})$: $x=\left\{{w_{s}^{t},h_{k}^{t}}\right\}$, and the witness $W$ is the updated local model $w_{k}^{t}$: $W=w_{k}^{t}$.
@@
    Each client $C_{k}$ transmits the proof $\pi_{k}$ for the updated local model’s performance metric and the hash of their updated local model $h_{k}^{t}$ to the server $S$, other than transmitting the updated local model $w_{k}^{t}$ to the server at this stage.
@@
    ${\text{cos}}\left(w_{k}^{t},w_{s}^{t}\right)=\frac{w_{k}^{t}\cdot w_{s}^{t}}{|w_{k}^{t}||w_{s}^{t}|}$ (1)

### v2

@@ IV-A Threat and Trust Model
  • Malicious clients may launch poisoning attacks by submitting falsified performance metrics of their local models, with the intent of disrupting the overall learning process.
@@ IV-B Veri-CS-FL
    Each client $C_{k}$ trains an updated local model $w_{k}^{t}$ from the last round global model $w^{t-1}$ utilizing its local dataset $D_{k}$, $k\in{1,2,\cdots,K}$ [^1] We assume that the local model training stage is conducted honestly by all clients, which is further guaranteed through a ZKP algorithm, as demonstrated in [7, 8, 9, 10, 11]. We only consider the model evaluation trust in the design of Veri-CS-FL. Subsequently, each client computes a performance metric of the updated local model $w_{k}^{t}$ by comparing the updated local model $w_{k}^{t}$ with the benchmark model $w_{s}^{t}$, and employs the proof generation algorithm of zk-SNRAK to generate a zero-knowledge proof $\pi_{k}$ for the performance metric of $w_{k}^{t}$: $\pi_{k}\leftarrow PROVE(PK,x,W)$, where the public input $x$ is the benchmark model of the server $w_{s}^{t}$ together with the hash of the updated local model $h_{k}^{t}=\text{hash}(w_{k}^{t})$: $x=\left\{{w_{s}^{t},h_{k}^{t}}\right\}$, and the witness $W$ is the updated local model $w_{k}^{t}$: $W=w_{k}^{t}$.
@@
    Each client $C_{k}$ transmits the proof $\pi_{k}$ for the updated local model’s performance metric, the hash of their updated local model $h_{k}^{t}$, and the corresponding performance metric to the server $S$, other than transmitting the updated local model $w_{k}^{t}$ to the server at this stage.
@@
    $\text{cos}_{k}^{t}\overset{\triangle}{=}{\text{cos}}\left(w_{k}^{t},w_{s}^{t}\right)=\frac{w_{k}^{t}\cdot w_{s}^{t}}{|w_{k}^{t}||w_{s}^{t}|}$ (1)